OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#17728 — Lyon optical loop
Scheduled Maintenance Report for Web Cloud
Completed
Following a performance drop of signal on the optical loop of Lyon, we will
work on the lal69 site to replace an attenuator.

Update(s):

Date: 2016-04-26 17:53:44 UTC
We have removed -5dB attenuator on the mpl69 site.

No more CRC are detected in the loop.

Date: 2016-04-26 17:51:37 UTC
Intervention complete. We have replaced attenuator 2dB by 10dB.

We will now work on mpl69.

Date: 2016-04-26 17:50:11 UTC
A technician is on site.
Posted Apr 26, 2016 - 17:49 UTC